OpenStudy (torialaw192):

Solve 4(x + 1) = 4x + 4 No solution 0 All real numbers 1

OpenStudy (torialaw192):

I'll give a medal to anyone that helps! ^^

pooja195 (pooja195):

Hi there! Welcome To OpenStudy! :) Any ideas on where to start?

OpenStudy (torialaw192):

Not really, I'm so lost on this one.

pooja195 (pooja195):

Are you familiar with the distributive property? \[\huge~\rm~\bf~a(b+c)=ab+ac \]

OpenStudy (torialaw192):

Yep

pooja195 (pooja195):

Perfect :~) 4(x + 1) = 4x + 4 So first let's start by distributing 4(x+1)=?

OpenStudy (torialaw192):

Put 4 into x and 1?

pooja195 (pooja195):

yep yep

OpenStudy (torialaw192):

Which will be 4x+4

pooja195 (pooja195):

Perfect so 4x+4=4x+4 Solve for x

OpenStudy (torialaw192):

Subrtract 4 from one side right?

pooja195 (pooja195):

yep

OpenStudy (torialaw192):

Which will be be 4x=4x+4-4 or 4x=4x

pooja195 (pooja195):

good keep going :)

OpenStudy (torialaw192):

And 4x=4x, divide both sides by four and x=x. So I'm thing it c. any solution.

pooja195 (pooja195):

Yep well done! :~)

OpenStudy (torialaw192):

Thank you! =3

pooja195 (pooja195):

You're welcome ^_^